OpenStudy (secret-ninja):

If f(x) = x - 4, what does f(3) represent? The value of (x - 4) when x = 3 The value of x when (x - 4) = 3 The value of 3f(x) when x = 3 The value of f(x) when (x - 4) = 3

OpenStudy (wmj259):

plug 3 in for x.

OpenStudy (solomonzelman):

I am going to give you a couple of examples. \(\large\color{black}{ f(x) = x - 4 }\) \(\large\color{blue}{ f(2) = (2) - 4 =-2 }\) \(\large\color{red}{ f(-3) = (-3) - 4 =-7 }\)

OpenStudy (secret-ninja):

f(3)=3-4 f(3)=-1 ?

OpenStudy (solomonzelman):

Yes

OpenStudy (secret-ninja):

So its A?

OpenStudy (wmj259):

f(x)=x+3 for example F(x) means when you input the value of x into the equation, where ever you see that variable x is where that number (in parenthesis goes).

OpenStudy (solomonzelman):

this is the value of f(x) when x=3. (Saying, in other words, the "f(3)", the f of 3")

OpenStudy (solomonzelman):

Yes it is A